OpenStudy (confuzion):

algebraically find the x-intercept(s) of this quadratic relation y= -x^2+2x+8

OpenStudy (3mar):

Could I help you?

OpenStudy (shadowlegendx):

You have to factor this quadratic, putting it into intercept form.

OpenStudy (shadowlegendx):

@confuzion

OpenStudy (confuzion):

How would i do that

OpenStudy (3mar):

"find the x-intercept(s) of this quadratic relation" means where does this parabola strike/intersect/touch the x-axis. In other words, it wants the roots (the solution) of this quadratic equation!

OpenStudy (confuzion):

I know the xintercepts are -2 and 4 but how do i algebraically find that out

OpenStudy (3mar):

Use the quadratic formula to find out these two, or as @ShadowLegendX has said; factor this quadratic!

OpenStudy (shadowlegendx):

\[y= -x^2+2x+8 \] Factor out -1 \[y= -(x^2 -2x-8)\] The product of 2 and -4 is -8 The sum of -4 and 2, is -2 \[y = -(x - 4)(x + 2)\]

OpenStudy (shadowlegendx):

Take the opposite sign of -4 and 2, Roots are 4 and -2

OpenStudy (confuzion):

so i don't do anything with the negative sign in front of the brackets

OpenStudy (shadowlegendx):

In the future, if you end up with parentheses that have something like (2x -5)(3x + 2) Solve for the roots by equaling them to zero, like so 2x - 5 =0 2x = 5 x = 5/2 3x + 2 = 0 3x = -2 x = -2/3

OpenStudy (shadowlegendx):

No, they just wanted the roots

OpenStudy (shadowlegendx):

or the "x-intercepts"

OpenStudy (confuzion):

okay thanks

OpenStudy (shadowlegendx):

Since they ask for the algebraic method, I would show your work how you got those roots x - 4 = 0 x = 4 x + 2 = 0 x = -2

OpenStudy (3mar):

\[\LARGE {y = -(x - 4)(x + 2)\\let~~ y=0\\-(x - 4)(x + 2)=0\\(x - 4)(x + 2)=0\\(x-4)=0~~~or~~~(x+2)=0\\x=4~~~~~~and ~~~~~~~x=-2}\]

OpenStudy (confuzion):

@ShadowLegendX AND @3mar Thanks

OpenStudy (3mar):

Don't mention it! That is with my pleasure! Any Help... Any Time... Any more questions?

OpenStudy (confuzion):

No but thank you so much!

OpenStudy (3mar):

Don't hesitate to ask me if there is any difficulty faces you!

OpenStudy (whpalmer4):

@confuzion the reason the negative sign in front of the factors doesn't matter is that you can pass an infinite variety of parabolas through a given pair of x-intercepts. Try graphing the following: \[y=-(x-1)(x+1)\]\[y=-2(x-1)(x+1)\]\[y=3(x-1)(x+1)\]on the same graph. You can do this by giving google the following search string: plot -1(x-1)(x+1), -2(x-1)(x+1), 3(x-1)(x+1)